OpenStudy (anonymous):

“Trish has a jug of contaminated water that she knows is only 75% clean. She also knows that the water will be safe to drink if she can mix it with 100% clean water so that it is 97% clean. How much clean water (100% clean) would Trish need to mix with the contaminated water in order to get 1 gallon of drinkable water?” A student used the following system of equations to model the situation given above. Which of the statements is true about the system of equations the student used?

OpenStudy (anonymous):

1) 75x+100y=97/(x+y) eq was incorrect 2) x+y=1 gallon=97/(75x+100y) eq was incorrect no guarantee